OpenStudy (anonymous):

Could someone please help me with this? A grease gun holds a cylindrical tube of grease that is 14 in. long and has a radius of 1.5 in. With each squeeze, 15 in3 of grease is used. How many times can the grease gun be squeezed before it is empty? Use 3.14 to approximate pi, and round your final answer to the nearest whole number.

OpenStudy (anonymous):

whats are ur answer choices ?

OpenStudy (anonymous):

as in what answers I've come up with or like as a test answer choices?

OpenStudy (anonymous):

test choices o_o

OpenStudy (anonymous):

o_o there's no choices with this question

OpenStudy (ribhu):

volume of gun = no of times it can be squeezed* 15

OpenStudy (anonymous):

so the volume, I got 31.5 when I found it though I don't know if thats correct, multiplied by 15 would be my answer?

OpenStudy (ribhu):

no its not

OpenStudy (ribhu):

3.14*2.25*14 = n *15 solve for n

OpenStudy (anonymous):

so 3.14 x 2.25 x 14 = 98.91 and then 98.91 x 15 = 1483.65

OpenStudy (ribhu):

noo--- it would be 98.91/15 = n

OpenStudy (anonymous):

oh, I thought it was multiplied because you had the * sign. So n would equal 6.594 but should that be rounded to the nearest whole number?

OpenStudy (ribhu):

yeah yo have to. 6 complete squeezes

OpenStudy (anonymous):

Okay, thank you for your help! c: